At a stop sign, some drivers come to a full stop, some come to a `rolling stop' (not a full stop, but slow down), and some do not stop at all. We would like to test if there is an association between gender and type of stop (full, rolling, or no stop). We collect data by standing a few feet from a stop sign and taking note of type of stop and the gender of the driver. Below is a contingency table summarizing the data we collected.MaleFemaleFull stop66Rolling stop1615No stop43

If gender is not associated with type of stop, how many males would we expect to not stop at all?



1
Expert's answer
2021-04-15T07:26:06-0400


Total = 50

If gender is not associated with type of stop, expectation to not stop at all is

= P(Male) "\times" (number of people that won’t stop)

"= \frac{26}{50} \times 7 \\

= 0.52 \times 7 \\

= 3.64"